OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − 2y = 2 3x + y = 6 Numerical Answers Expected! I just need the answer in a hurry!!!

OpenStudy (anonymous):

@Hero @jose12345 @dillon754 @deshawn1 @quickstudent @oliviadetataaa @popitree @kaylaprincess @HatcrewS @bianca. @bg45

OpenStudy (anonymous):

x=2+2y from 1st eq. put value of x into 2nd eq. 3(2+2y)+y=6 6+6y+y=6 7y=0 y=0 and x=2+2.0=2

OpenStudy (anonymous):

ty